# VioPTT: Violin Technique-Aware Transcription from Synthetic Data Augmentation

### v2

@@ Abstract
While automatic music transcription is well-established in music information retrieval, most models are limited to transcribing pitch and timing information from audio, and thus omit crucial expressive and instrument-specific nuances. One example is playing technique on the violin, which affords its distinct palette of timbres for maximal emotional impact. Here, we propose VioPTT (Violin Playing Technique-aware Transcription), a lightweight, end-to-end model that directly transcribes violin playing technique in addition to pitch onset and offset. Furthermore, we release MOSA-VPT, a novel, high-quality synthetic violin playing technique dataset to circumvent the need for manually labeled annotations. Leveraging this dataset, our model demonstrated strong generalization to real-world note-level violin technique recordings in addition to achieving state-of-the-art transcription performance. To our knowledge, VioPTT is the first to jointly combine violin transcription and playing technique prediction within a unified framework.
@@ 1 Introduction
The goal of automatic music transcription (AMT) is to convert an audio recording into a symbolic representation containing pitch and timing information. As one of the most studied tasks in music information retrieval (MIR), AMT has achieved substantial progress with the advent of deep learning. In particular, instrument-specific models have been developed for piano [1, 2, 3, 4], guitar [5, 6], and violin [7], while more recent approaches such as BasicPitch [8] and MT3 [9] provide instrument-agnostic solutions. Despite these advances, most AMT systems are limited to note transcription and overlook expressive, instrument-specific performance nuances.
@@
To overcome these challenges, we propose a lightweight end-to-end violin transcription model that extends conventional AMT by jointly predicting pitch, onset, offset and playing technique. We show that state-of-the-art transcription performance is achievable by training from well-aligned solo-violin performances with augmentation and without relying on pretrained representations from other instruments. We further introduce and release a synthetic violin playing technique dataset to eliminate the need for expert technique annotations. Despite training on synthetic data, our model is able to successfully classify playing technique in real-world violin audio. Our approach thus provides a richer and more expressive representation of violin performance, and expand AMT towards capturing the full depth of musical expression.
@@ 2 Related Work
Regarding violin technique classification, earlier work focused on modeling physical attributes of bow motion across different techniques [14]. Subsequent developments by Su et al. [15] demonstrated the feasibility of classifying playing technique in single notes using a support vector machine with sparse-coded magnitude and phase-derivative features from extracted from audio. Kruger et al. [16] later expanded the model to allow simultaneous technique classification across violin, viola, cello, and double bass using augmented audio features extracted from a Hartley transform. A CNN-based violin technique classifier on single notes has also been recently proposed in [17].
@@ 3.1 Playing Technique-Aware Transcription Model
We adopt Kong et al. [1]’s high-resolution piano model for violin audio in the transcription module. One CRNN module is trained to separately predict at the frame-level onset, offset, velocity as regression targets, and frame as a binary target. Each CRNN comprises four convolutional blocks followed by two bidirectional GRU layers. Features are then passed through a fully connected layer, two biGRUs, and a final 88-dimension sigmoid layer covering the violin pitch range.
@@ 3.2 Data augmentation
3.2.1 Pitch and Timing Augmentation
@@ 4.1 Datasets
MOSA-VPT. Using the synthesis pipeline described in Section 3.2.2, we construct MOSA-VPT, a large-scale synthetic dataset of violin playing techniques111link to dataset will be made available upon acceptance. This dataset comprises 76 hours of audio-MIDI pairs balanced across four techniques: détaché, flageolet, spiccato, and pizzicato. We target these four techniques because they exhibit distinct acoustic signatures and cannot be trivially modeled by existing transcription approaches (e.g., tremolo may be approximated as rapid détaché). This corpus provides note- and technique-aligned supervision at scale and is used to train our technique-aware transcription model.
@@ 4.2 Implementation
For pitch and timing transcription, performance was assessed using Precision (P), Recall (R), F1-score (F1), and onset-only F1-score (F1) in the mir_eval library [25]. For P, R, and F1, a note is regarded as correct if its pitch deviation is within 50 cents, onset deviation is within 50 ms, and offset deviation is within 20% of the note duration. For F1, the same criteria for F1 is applied without considering offsets. We used the existing state-of-the-art violin transcription model MUSC [7] as baseline. For technique transcription, performance was evaluated in two ways: First, using macro accuracy, defined as the mean of accuracy values computed independently for each class to account for label imbalance; Second, using per-class accuracy.
@@ 5.2 Playing Technique Transcription
Table 2 presents the violin technique transcription results on the RWC dataset under different transcription feature ablations in the articulation module. In the No ablation condition, the model achieved the highest macro accuracy of 77.22%, confirming the benefit of having all transcription information. Removing individual features degraded overall performance: excluding Offset led to a large drop to 59.71%, while excluding Frame or Onset also reduced accuracy to a smaller extent.
Examining each technique individually, flageolet achieved the highest accuracy when onset information was removed (91.55%), followed by full ablation. This suggests that the model primarily relied on harmonic features to classify flageolet. For détaché, accuracy improved when frame features were excluded (70.16% vs. 63.12% without ablation). This indicates that binary frame information may have introduced noise, whereas offset and onset cues seemed to be more discriminative. Pizzicato showed strongest performance with no ablation (88.80%), and excluding frame or onset only moderately reduced accuracy. However, excluding velocity proved to be detrimental and resulted in near-zero accuracy. This suggests that the model may have exploited velocity as a defining characteristic for this technique. Spiccato accuracy was similar without or with full ablation, with substantial performance drops when timing information was ablated. These indicate that classification relied on the combination of cues–especially timing information. Overall, the ablation study show that violin techniques exhibit distinct dependencies on transcription features: frame- and onset-based cues were particularly critical for techniques with sharp articulations such as détaché and spiccato, while harmonic cues dominated flageolet.
@@ 6 Conclusion and Future Work
While pitch and timing information remain fundamental in AMT, transcribing expressive nuances from different playing techniques is a crucial next step towards fine-grained music performance modeling. Here, we presented VioPTT, a light-weight end-to-end technique-aware transcription model for violin, and release MOSA-VPT, a high-quality synthetic violin playing technique dataset. In addition to exceeding the state-of-the-art in violin pitch and timing transcription, our model demonstrated high accuracy in transcribing playing technique in real-life performances despite training only from labeled synthetic data. This demonstrates the feasibility of using synthetic data as a proxy to technique annotations beyond single notes to obtain a richer and more nuanced transcription of expressive music performance. We envision our work to see applications in synthesis, performance analysis, and pedagogy. For future work, we aim to capitalize on our data synthesis approach to include more a wider variety techniques, as well as to extend the model to other bowed string instruments.

### v3

@@ Abstract
While automatic music transcription is well-established in music information retrieval, most models are limited to transcribing pitch and timing information from audio, and thus omit crucial expressive and instrument-specific nuances. One example is playing technique on the violin, which affords its distinct palette of timbres for maximal emotional impact. Here, we propose VioPTT (Violin Playing Technique-aware Transcription), a lightweight cascade model that directly transcribes violin playing technique in addition to pitch onset and offset. Furthermore, we release MOSA-VPT, a novel, high-quality synthetic violin playing technique dataset to circumvent the need for manually labeled annotations. Leveraging this dataset, our model demonstrated strong generalization to real-world note-level violin technique recordings in addition to achieving state-of-the-art transcription performance. To our knowledge, VioPTT is the first to jointly combine violin transcription and playing technique prediction within a unified framework.
@@ 1 Introduction
The goal of automatic music transcription (AMT) is to convert an audio recording into a symbolic representation for music analysis and performance understanding. As one of the most studied tasks in music information retrieval, AMT has achieved substantial progress with the advent of deep learning. In particular, instrument-specific models have been developed for piano [12, 22, 6, 7], guitar [9, 20], and violin [21], while more recent approaches such as BasicPitch [1] and MT3 [4] provide instrument-agnostic solutions. Despite these advances, most AMT systems are limited to note transcription and overlook expressive, instrument-specific performance nuances.
@@
To overcome these challenges, we propose a lightweight cascade violin transcription model that extends conventional AMT by jointly predicting pitch, onset, offset and playing technique. We show that state-of-the-art transcription performance is achievable by training from well-aligned solo-violin performances with augmentation and without relying on pretrained representations from other instruments. We further introduce and release a synthetic violin playing technique dataset to eliminate the need for expert technique annotations. Despite training on synthetic data, our model is able to successfully classify playing technique in real-world violin audio. Our approach thus provides a richer and more expressive representation of violin performance, and expand AMT towards capturing the full depth of musical expression. For reproducibility, model and code are available at https://github.com/y10ab1/VioPTT.
@@ 2 Related Work
Recent work has also explored frame-level playing technique detection beyond note-based formulations. Li et al. [14, 16] proposed a frame-level multi-label approach with multi-scale convolution and self-attention to handle overlapping and mixed techniques in continuous performances. Building on this formulation, MERTech [15] further incorporated self-supervised music representations and multi-task finetuning with auxiliary pitch and technique-onset prediction, achieving state-of-the-art results across multiple instrument technique datasets.
@@ 3.1 Playing Technique-Aware Transcription Model
We adopt Kong et al. [12]’s high-resolution piano model for violin audio in the transcription module. One CRNN module is trained to separately predict at the frame-level onset, offset, velocity as regression targets, and note activation as a binary target. Each CRNN comprises four convolutional blocks followed by two bidirectional GRU layers. Features are then passed through a fully connected layer, two biGRUs, and a final 88-dimension sigmoid layer covering the violin pitch range.
@@ 3.2 Data augmentation
3.2.1 Semitone-Level Pitch and Timing Transcription
@@ 4.1 Datasets
MOSA-VPT. Using the synthesis pipeline described in Section 3.2.2, we construct MOSA-VPT, a large-scale synthetic dataset of violin playing techniques111Available at https://doi.org/10.5281/zenodo.18295471. This dataset comprises 76 hours of audio-MIDI pairs balanced across four techniques: détaché, flageolet, spiccato, and pizzicato. We target these four techniques because they exhibit distinct acoustic signatures and cannot be trivially modeled by existing transcription approaches (e.g., tremolo may be approximated as rapid détaché). This corpus provides note- and technique-aligned supervision at scale and is used to train our technique-aware transcription model.
@@ 4.2 Implementation
For pitch and timing transcription, performance was assessed using Precision (P), Recall (R), F1-score (F1), and onset-only F1-score (F1) in the mir_eval library [19]. For P, R, and F1, a note is regarded as correct if its pitch deviation is within 50 cents, onset deviation is within 50 ms, and offset deviation is within 20% of the note duration. For F1, the same criteria for F1 is applied without considering offsets. For technique transcription, performance was evaluated in two ways: First, using macro accuracy, defined as the mean of accuracy values computed independently for each class to account for label imbalance; Second, using per-class accuracy.
As baselines, we compared our work against the existing state-of-the-art violin transcription model MUSC [21] and MERTech [15], a state-of-the-art multi-label technique classification model. As MERTech relies on technique-annotated data, it was trained exclusively on the MOSA-VPT dataset following the original experimental setting for 200 epochs. Although MERTech is not explicitly designed for note transcription, pitch and onset information are incorporated as auxiliary cues and so we evaluated MERTech on both note and technique transcription.
@@ 5.2 Playing Technique Transcription
Table 2 presents the violin technique transcription results on the RWC dataset under different transcription feature ablations in the articulation module. With the no ablation condition, our model achieved the highest macro accuracy of 77.22%, confirming the benefit of having all transcription information. Removing individual features degraded overall performance: excluding offset led to a large drop to 59.71%, while excluding frame or onset also reduced accuracy to a smaller extent.
Examining each technique individually, MERTech [15] achieved the highest accuracy for flageolet, with VioPTT obtaining the second-best performance when onset information was excluded. This suggests that our model primarily relied on harmonic features for flageolet classification. For détaché, accuracy improved when frame features were excluded. This indicates that binary frame information may have introduced noise, whereas offset and onset cues seemed to be more discriminative. Pizzicato showed strongest performance with no ablation, and excluding frame or onset only moderately reduced accuracy. However, excluding velocity proved to be detrimental and resulted in near-zero accuracy. This suggests that the model may have exploited velocity as a defining characteristic for this technique. Spiccato accuracy was similar without or with full ablation, with substantial performance drops when timing information was ablated. These indicate that classification relied on the combination of cues–especially timing information. Overall, the ablation study show that violin techniques exhibit distinct dependencies on transcription features: frame- and onset-based cues were particularly critical for techniques with sharp articulations such as détaché and spiccato, while harmonic cues dominated flageolet.
@@ 6 Conclusion and Future Work
While pitch and timing remain fundamental in AMT, modeling expressive nuances across playing techniques is essential for fine-grained performance transcription. We present VioPTT, a lightweight, technique-aware transcription model for violin, together with MOSA-VPT, a high-quality synthetic dataset for violin playing techniques. Our model achieves state-of-the-art performance in pitch and timing transcription and accurately recognizes playing techniques in real-world performances despite being trained only on synthetic data. This demonstrates the feasibility of using synthetic data as a proxy to technique annotations beyond single notes to obtain a richer and more nuanced transcription of expressive music performance. Future work will extend our synthesis framework to a broader range of techniques and other bowed string instruments.